A system has 4 input lines {A,B,C,D} and 1 output line {F}. The input lines {A, B, C} of the system are connected 

to the selection lines {S2 , S1 , S0} of an 8×1 multiplexer, respectively. The connection configuration of data 

input lines {i0, i1, … , i7 } of the multiplexer are given as follows: 

i1= i2 = D; i3 = i5 =1; i0 =i6 = D' ; and i7 = i4 = 0 

Determine the Boolean function that the output of the multiplexer implements. Also, provide a minimized SoP 

expression for the obtained Boolean function F(A,B,C,D).


1
Expert's answer
2021-05-26T15:18:02-0400

(a) Here GT A,B,C, = "S_2,S_1,S_0" respectively

G.T "I_1=I_2=D" "I_3=I_5=D" "I_7=I_4=0"

"F(A,B,C,D)=\bar{A}\bar{B}CD+\bar{A}B\bar{C}D+\bar{A}BC\bar{D}+A\bar{B}C\bar{D}"



Minimized = "\sum f (3,5,6,12)"

sop form

f(A,,B,C,D)

(b) "I_1=I_4=D" "I_3=I_5=1" "I_0=I_6=B_0"

"I_7=I_2=D"

substitute in the above expression ,

F(A,B,C,D)= "\bar{A}\bar{B}CD+\bar{A}\bar{B}\bar{C}\bar{D}+\bar{A}BC+A\bar{B}\bar{C}+A\bar{B}\bar{C}\bar{D}+A\bar{B}C+ABC\bar{D}"




Y="\bar{S_2,S_1,S_0} I_0+S_2S_1S_0I_1+S_2S_1S_0I_2+S_2S_1S_0I_3"

"S_2S_1S_0I_4+S_2S_1S_0I_5+S_2S_1S_0I_6+S_2S_1S_0I_7"

"F(A,B,C,D)=\bar{A}\bar{B}CD+\bar{A}\bar{B}\bar{C}\bar{D}+\bar{A}BC+A\bar{B}\bar{C}+A\bar{B}\bar{C}\bar{D}+A\bar{B}C+ABC\bar{D}"

"\bar{A}BC=\bar{A}BC(D+\bar{D})" as "D+\bar{D}=1"

"F(A,B,C,D)=\sum(3,4,6,7,8,10,11,13)"

minimal pos expression

"\pi(0,1,2,5,9,12,14,15)"


(c) F(A,B,C,D)= CD+ABCD+ABC+ABCD

"I" is connected to "I_7" D s connected to "I_1"

D is connected to "I_5,I_0,I_2,I_4,1_6" 0 a connected to "I_3"